OpenStudy (anonymous):

A box contains 4 red, 5 green, and 6 blue balls. You take two balls at random without replacement. Find the probability that both balls are red.

OpenStudy (anonymous):

Total balls = 6+4+5 = 15 Red balls = 4 First red ball = 4/15 Second red ball = 3/14 total probability = First X Second P = 4/15 * 3/14